OpenStudy (b2st):

._. I don't understand how to get the answer. lol Proglem.. find f of g f(x)= (3 root) x-1 Uhh the three is front of the root sign and the x-1 is under the root. I'll retype it here one sec... Oh the answer is apparently x..

OpenStudy (b2st):

\[\sqrt[3]{x-1}\]

OpenStudy (b2st):

^ this what I meant. lol and the answer is apparently x. How do you get that?

OpenStudy (anonymous):

you could say \(g(x)=x-1\) and \(f(x)=\sqrt[3]{x}\) which would make \[f\circ g(x)=f(x-1)=\sqrt[3]{x-1}\]

OpenStudy (b2st):

Thanks! But I'm still confuse. o.o Why did you do that?

OpenStudy (anonymous):

because that is what you asked right? write \[\sqrt[3]{x-1}\] as a composite function

OpenStudy (b2st):

Yes, I did. My math teacher didn't really lecture about this part of solving. O ^ O

OpenStudy (anonymous):

what would you do if you had a number for \(x\)? like say \(x=9\) 1) subtract 1 to get \(9-1=8\)(that is the first function) 2) take the cubed root (that is the second function)